# Decentralized Stochastic Nonconvex Optimization under the Relaxed Smoothness

### v1

@@ 5 Discussion and Related Work
The design of our DSNGD (Algorithm 2) is based on normalized (stochastic) gradient descent, which is extensively studied in centralized scenario [34, 4, 9, 15, 11]. Recently, Li et al. [25] studied normalized stochastic gradient descent in decentralized setting, while their analysis is only limited to the setting of standard smoothness. In a very recent work, Jiang et al. [17, Algorithm 2 and Theorem 4] claims a first-order method for decentralized optimization under the relaxed smoothness, while each of its iteration requires every node to perform the step of gradient clipping which depends on accessing the mean vector ${\bar{\bf{x}}}_{i}^{t}$ and the corresponding exact local gradient $\nabla f_{i}({\bar{\bf{x}}}^{t})$. Recall that the decentralized method typically only allows each node directly communicates with its neighbors (Assumption 5) and the stochastic first-order oracle can only access the unbiased gradient estimator with bounded variance (Assumption 4). Therefore, we argue that their algorithm is not truly decentralized or stochastic. In contrast, our method only requires each node to iterate with its local stochastic gradient (line 9 of Algorithm 2), which is fully decentralized and stochastic.
The convergence analysis for our DSNGD (Algorithm 2) depends on the assumption of bounded gradient dissimilarity (Assumption 3), which is commonly used in distributed optimization [47, 13, 2, 26]. Specifically, we use this assumption to establish the inequality $\left\|\nabla f({\bar{\bf{x}}}^{t+1})\right\|\leq(L_{0}+L_{1}\zeta)\eta+\left(1+L_{1}\eta\right)\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|$, which connects the gradients at points gradients at points ${\bar{\bf{x}}}^{t}$ and ${\bar{\bf{x}}}^{t+1}$ (see Lemma 9 in Appendix C). We also combine Assumption 3 with the $(L_{0},L_{1})$-smoothness of local functions (Assumption 2) to guarantee the global objective $f=\frac{1}{m}\sum_{i=1}^{m}f_{i}$ is $(L_{0}+L_{1}\zeta,L_{1})$-smooth (Appendix B). How to establish the convergence for decentralized relaxed smooth optimization without Assumption 3 is still an open problem. It is worth noting that simply assuming that each local function $f_{i}$ is $(L_{0},L_{1})$-smooth even cannot guarantee that the global function $f=\frac{1}{m}\sum_{i=1}^{m}f_{i}$ is also $(L_{0},L_{1})$-smooth, which is different with the standard smooth case. We illustrate this point with the following example.
@@
Recall that the discussion after Corollary 1 shows that the sample complexity of our method matches the lower bound in the special case of $L_{1}=0$ [32]. This implies that the ${\mathcal{O}}(\epsilon^{-4})$ dependence on accuracy we obtained cannot be improved in our setting. If we target to achieve an $\mathcal{O}(\epsilon^{-3})$ dependence in the sample complexity, the assumption of expected relaxed smoothness is necessary [4, 40], e.g., strengthen condition (1) in Assumption 2 to ${\mathbb{E}}[\left\|\nabla F_{i}({\bf{x}};{\bm{\xi}}_{i})-\nabla F_{i}({\bf{y}};{\bm{\xi}}_{i})\right\|^{2}]\leq\left(L_{0}+L_{1}\left\|\nabla f_{i}({\bf{x}})\right\|\right)^{2}\left\|{\bf{x}}-{\bf{y}}\right\|^{2}$, then applying the variance-reduced gradient estimators [4, 40] on local agents is possible to achieve the sample complexity depends on $\mathcal{O}(\epsilon^{-3})$.
@@ 6 Experiments
In this section, we empirically compare our decentralized normalized stochastic gradient descent (DNSGD) with baseline methods, including decentralized stochastic gradient descent (D-SGD) [26], decentralized stochastic gradient descent with gradient tracking (D-SGT) [36, 39], and decentralized normalized averaged stochastic approximation (D-NASA) [25]. The experiments are performed on a virtualized cluster with 8 vGPUs (NVIDIA RTX 5090, 25.6 GB memory), 8 vCPUs (Intel Xeon Gold 6530), and 96 GB memory.
We present the numerical results on the application of image classification for the MNIST dataset. Specifically we train a three-layer neural network, i.e., the first layer consists of a convolutional layer with one input channel and $16$ output channels, followed by batch normalization, a ReLU activation function, and max pooling; the second layer mirrors the first but has $16$ input channels and $32$ output channels; the last layer is a fully connected layer that predicts $10$ classes.
We conduct the experiments on a distributed system with $10$ agents and its topology is based on the Erdös-Rényi graph [10] with the connectivity probability of $p=0.6$. For all algorithms, we set batch size to $b=256$. The step size of D-SGD, D-SGT and DNSGD are chosen from $\{0.0001,0.0005,0.001,0.005,0.01,0.05,0.1,0.5,1\}$. For our DNSGD, we set the iterations number of Chebyshev acceleration (Algorithm 1) to $\hat{K}=K=2$. We show the results for the gradient norm and the test accuracy in Figures 1 and 2, demonstrating that the proposed DNSGD outperforms all baseline methods.
Refer to caption Refer to caption
(a) sample complexity vs. gradient norm (b) communication rounds vs. gradient norm
Figure 1: We present the sample complexity and the communication rounds against the average of gradient norms $\frac{1}{m}\sum_{i=1}^{m}\nabla f_{i}({\bf{x}}_{i}^{t})$ for MNIST dataset.
Refer to caption Refer to caption
(a) sample complexity vs. test accuracy (b) communication rounds vs. test accuracy
@@
Figure 2: We present the sample complexity and the communication rounds against the test accuracy for MNIST dataset.
@@ 7 Conclusion
This paper proposes decentralized normalized stochastic gradient descent for nonconvex optimization under the relaxed smoothness. We provide theoretical analysis to show our method can achieve the $\epsilon$-stationary point on each agent. The empirical studies show that our method performs better than baselines. In future work, we would like to establish decentralized methods for convex optimization under the relaxed smoothness [5, 12, 14, 20, 31, 46]. It is also interesting to extend our results to the more general assumption of symmetric relaxed smoothness [4] and the stochastic gradient estimator with heavily tail-noise [30].
@@ Appendix A Some Technical Lemmas
This result is presented in the analysis Li et al. [25, Lemma B.5]. Here, we provide the detailed proof for completeness. We have
@@
This result can be obtained by following Luo et al. [33, Lemma 2]. Here, we provide the detailed proof for completeness. The update in line 4 and Proposition 2 imply ${\bar{\bf{v}}}^{0}={\bar{\bf{g}}}^{0}$. Suppose it holds ${\bar{\bf{v}}}^{t}={\bar{\bf{g}}}^{t}$, then the update in line 10 means
@@
Lemma 7 (Zhang et al. [53, Appendix A.1]).
@@
Proposition 1 means the objective $f$ is $(L_{f},L_{1})$-smooth, then applying equation (29) of Zhang et al. [53, Appendix A.1] directly achieves equation (7). ∎
@@ Appendix H The Proof of Proposition 3
is increasing. Taking z=yxz=y-x, we have

### v2

@@ 5 Discussion and Related Work
The design of our DNSGD (Algorithm 2) is based on normalized (stochastic) gradient descent, which is extensively studied in centralized scenario [35, 4, 9, 15, 11]. In a related work, [26] studied normalized stochastic gradient descent in decentralized setting, while their analysis is only limited to the setting of standard smoothness. In a very recent work, Jiang et al. [17, Algorithm 2 and Theorem 4] claims a first-order method for decentralized optimization under the relaxed smoothness, while each of their iteration requires every node to perform the step of gradient clipping which depends on accessing the mean vector ${\bar{\bf{x}}}_{i}^{t}$ and the corresponding exact local gradient $\nabla f_{i}({\bar{\bf{x}}}^{t})$. Recall that the decentralized method typically only allows each node directly communicates with its neighbors (Assumption 5) and the stochastic first-order oracle can only access the unbiased gradient estimator with bounded variance (Assumption 4). Therefore, we argue that their algorithm is not truly decentralized or stochastic. In contrast, our method only requires each node to iterate with its local stochastic gradient (line 9 of Algorithm 2), which is fully decentralized and stochastic.
The convergence analysis for our DNSGD (Algorithm 2) depends on the assumption of bounded gradient dissimilarity (Assumption 3), which is commonly used in distributed optimization [48, 13, 2, 27]. Specifically, we use this assumption to establish the inequality $\left\|\nabla f({\bar{\bf{x}}}^{t+1})\right\|\leq(L_{0}+L_{1}\zeta)\eta+\left(1+L_{1}\eta\right)\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|$, which connects the gradients at the points ${\bar{\bf{x}}}^{t}$ and ${\bar{\bf{x}}}^{t+1}$ (see Lemma 9 in Appendix C). We also combine Assumption 3 with the $(L_{0},L_{1})$-smoothness of local functions (Assumption 2) to guarantee the global objective $f$ is $(L_{0}+L_{1}\zeta,L_{1})$-smooth (see the proof in Appendix B). However, how to establish the convergence for decentralized relaxed smooth optimization without bounded gradient dissimilarity (Assumption 3) is still an open problem. It is worth noting that simply assuming that each local function $f_{i}$ is $(L_{0},L_{1})$-smooth even cannot guarantee that the global function $f=\frac{1}{m}\sum_{i=1}^{m}f_{i}$ is also $(L_{0},L_{1})$-smooth, which is different with the standard smooth case. We illustrate this point with the following example.
@@
Recall that the discussion after Corollary 1 shows that the sample complexity of our method matches the lower bound in the special case of $L_{1}=0$ [33]. This implies that the ${\mathcal{O}}(\epsilon^{-4})$ dependence on accuracy we obtained cannot be improved in our setting. If we target to achieve the sharper $\mathcal{O}(\epsilon^{-3})$ dependence in the sample complexity, the assumption of expected relaxed smoothness is necessary [4, 41], e.g., strengthen condition (1) in Assumption 2 to ${\mathbb{E}}[\left\|\nabla F_{i}({\bf{x}};{\bm{\xi}}_{i})-\nabla F_{i}({\bf{y}};{\bm{\xi}}_{i})\right\|^{2}]\leq\left(L_{0}+L_{1}\left\|\nabla f_{i}({\bf{x}})\right\|\right)^{2}\left\|{\bf{x}}-{\bf{y}}\right\|^{2}$, then applying the variance-reduced gradient estimators [4, 41, 11] at local agents is possible to achieve the sample complexity depends on $\mathcal{O}(\epsilon^{-3})$.
@@ 6 Experiments
In this section, we empirically compare our decentralized normalized stochastic gradient descent (DNSGD) with baseline methods, including decentralized stochastic gradient descent (D-SGD) [27], decentralized stochastic gradient descent with gradient tracking (D-SGT) [37, 40], and decentralized normalized averaged stochastic approximation (D-NASA) [26]. The experiments are performed on a cluster with 8 virtual GPUs, 8 virtual CPUs, and 96 GB memory.
We present numerical results on the applications of image classification for MNIST and Fashion-MNIST datasets. The MNIST dataset is a well-known collection of handwritten digits (10 classes), containing $70,000$ images, with $60,000$ for training and $10,000$ for testing [23]. The Fashion-MNIST dataset serves as a more challenging variant, consisting of $70,000$ images of fashion items with 10 classes, such as trouser and dress [49]. Each example of MNIST or Fashion-MNIST is a $\times28$ grayscale image.
Refer to caption  Refer to caption Refer to caption
(a) sample vs. accuracy (b) communication vs. accuracy (c) linear speed-up
@@
Figure 1: We present numerical results on the MNIST dataset over the Erdös-Rényi network. The subfigures (a) and (b) present the sample complexity per agent and the communication rounds against the average test accuracy on the network with $m=16$ agents. The subfigure (c) verifies the linear speedup effect of our DNSGD on the network with $m=2,4,8,16$ agents.
Refer to caption  Refer to caption Refer to caption
(a) sample vs. accuracy (b) communication vs. accuracy (c) linear speed-up
Figure 2: We present numerical results on the Fashion-MNIST dataset over the Erdös-Rényi network. The subfigures (a) and (b) present the sample complexity per agent and the communication rounds against the average test accuracy on the network with $m=16$ agents. The subfigure (c) verifies the linear speedup effect of our DNSGD on the network with $m=2,4,8,16$ agents.
Our model employs a four-layer neural network, consisting of two convolutional layers and two fully connected layers. The first convolutional layer processes the single-channel input image with $32$ kernels of size $3\times 3$, followed by ReLU activation and $2\times 2$ max-pooling. The second convolutional layer takes the $32$-channel feature maps from the first layer, applies another $32$ kernels of size $3\times 3$, and is followed by ReLU activation and $2\times 2$ max-pooling. To enhance generalization, we apply dropout with a rate of $0.25$ after the convolutional layers. The resulting feature maps are then flattened and passed through a fully connected layer with $128$ units and ReLU activation, followed by a second dropout layer with a rate of $0.5$. Finally, we use a fully connected output layer to generate scores for the $10$ classes.
We perform numerical comparisons on a distributed system with $16$ agents, where the data is evenly distributed among them. In our experiments, we employ two network topologies, i.e., the Erdös-Rényi graph [10] with the connectivity probability of $p=0.6$ and the ring graph. For all algorithms, we set batch size to $b=250$. The step size of D-SGD, D-SGT and DNSGD are chosen from $\{0.0001,0.0005,0.001,0.005,0.01,0.05,0.1,0.5\}$. The learning rate of the D-NASA algorithm is given by $\eta_{t}=m^{1/4}t^{3/4}$ which follows the theory and implementation of Li et al. [26], where $t$ is the iteration counter and $m$ is the number of agents. For our DNSGD, we set the iterations number of Chebyshev acceleration (Algorithm 1) to $\hat{K}=K=2$. We present the numerical results on the Erdös-Rényi graph in Figures 1 and 2, and the results on the ring graph are shown in Figures 3 and 4. We evaluation the performance by showing the sample complexity and the communication rounds against the average test accuracy, i.e., the mean of the test accuracies achieved by the models from all agents. We can observe that DNSGD outperforms all baseline methods. The linear speed-up of our method is verified by comparing the results on the networks with $m=2,4,8,16$ agents. Figures 1(c), 2(c), 3(c), and 4(c) show that the more agents do lead to the faster convergence.
Refer to caption  Refer to caption Refer to caption
(a) sample vs. accuracy (b) communication vs. accuracy (c) linear speed-up
Figure 3: We present numerical results on the MNIST dataset over the ring network. The subfigures (a) and (b) present the sample complexity per agent and the communication rounds against the average test accuracy on the network with $m=16$ agents. The subfigure (c) verifies the linear speedup effect of our DNSGD on the network with $m=2,4,8,16$ agents.
Refer to caption  Refer to caption Refer to caption
(a) sample vs. accuracy (b) communication vs. accuracy (c) linear speed-up
Figure 4: We present numerical results on the Fashion-MNIST dataset over the ring network. The subfigures (a) and (b) present the sample complexity per agent and the communication rounds against the average test accuracy on the network with $m=16$ agents. The subfigure (c) verifies the linear speedup effect of our DNSGD on the network with $m=2,4,8,16$ agents.
@@ 7 Conclusion
This paper proposes decentralized normalized stochastic gradient descent for nonconvex optimization under the relaxed smoothness. We provide theoretical analysis to show our method can achieve the $\epsilon$-stationary point on each agent. In particular, we design a new Lyapunov function including the product of the gradient norm and the consensus error for convergence analysis, which successfully address the non-Lipschitz and unbounded local gradients in relaxed smooth decentralized optimization. The empirical studies also show that our method performs better than baselines. In future work, we would like to establish decentralized methods for convex optimization under the relaxed smoothness [5, 12, 14, 20, 32, 47]. It is also interesting to extend our results to the more general settings such as the symmetric relaxed smoothness [4] and the stochastic gradient estimator with heavy-tailed noise [31].
@@ Appendix A Some Technical Lemmas
This result is presented in the analysis Li et al. [26, Lemma B.5]. Here, we provide the detailed proof for completeness. We have
@@
This result can be obtained by following Luo et al. [34, Lemma 2]. Here, we provide the detailed proof for completeness. The update in line 4 and Proposition 2 imply ${\bar{\bf{v}}}^{0}={\bar{\bf{g}}}^{0}$. Suppose it holds ${\bar{\bf{v}}}^{t}={\bar{\bf{g}}}^{t}$, then the update in line 10 means
@@
Lemma 7 (Zhang et al. [55, Appendix A.1]).
@@
Proposition 1 means the objective $f$ is $(L_{f},L_{1})$-smooth, then applying equation (29) of Zhang et al. [55, Appendix A.1] directly achieves equation (7). ∎
@@ Appendix H The Proof of Proposition 3
is increasing. Taking $z=y-x\neq 0$, we have